The amount of sweetener used in a liter of milk tea is a normally distributed with a mean of 0.75 cup and a standard deviation of 0.6 cup. 

a. Suppose 10 liters of milk tea are randomly selected, what is the probability that it will have at least a cup of sweetener?

b.Suppose 20 liters of milk tea are randomly selected, what is the probability that it will have less than half a cup of sweetener?

c.15 liters of milk tea is randomly selected, what is the probability that it will have 0.50 to 1.25 cups of sweetener?


1
Expert's answer
2022-03-08T10:54:01-0500

Mean ("\mu") = 0.75

Standard deviation ("\u03c3") = 0.6

(a) at least 1 Cup


"Z=\dfrac{X-\mu}{σ}"


"Z=\dfrac{1-0.75}{0.6}"


"= 0.4166"

From the standard normal distribution tables

"Z(1)=0.6591"

Probability of at least 1 Cup

"P=1-0.6591 = 0.3409"


(b) Less than 0.5 cup

"Z=\dfrac{X-\mu}{σ}"


"Z=\dfrac{0.5-0.75}{0.6}"


"Z=-0.4166"


From the standard normal distribution tables probability of less than 0.5 cups

"P(0.5)=0.3409"


(c) Between 0.5 and 1.25


"Z(1.25)=\dfrac{1.25-0.75}{0.6}=0.8333"


"Z(0.5)=-0.4166"


From the standard normal distribution tables probability of between 0.5 and 1.25 cups is:


"=1-[P(0.5)+P(1.25)]"


"=1-[0.3409+0.20327]"


"=0.45583"